# thanks: These authors co-supervised the projectthanks: These authors co-supervised the project

### v1

@@ Abstract
Dissipation engineering offers a powerful tool for quantum technologies. Recently, new superconducting devices demonstrated an engineered two-photon dissipation rate exceeding all other relevant timescales. In particular, they have proven most useful to prevent transitions between the logical states |±$|\pm\alpha\rangle$ of a cat qubit. Here, we present three key applications of strong two-photon dissipation for quantum measurement and control, beyond cat qubit stabilization. Firstly, we demonstrate its efficacy in overcoming limitations encountered in Wigner tomography at high photon numbers. Secondly, we showcase its potential for realizing universal gates on cat qubits, exploiting the coherent mapping between cat qubit states and superpositions of 0 and 1 photons. Finally, we harness the transient dynamics of a cat state under two-photon dissipation to prepare squeezed cat states with a squeezing factor exceeding 3.8 dB.
@@
Existing devices [9, 11, 12, 16, 18, 19] that demonstrate two-photon dissipation use a dedicated mode, denoted as the buffer, to mediate the interaction between the memory and its environment. In our experiment, this buffer dissipates photons at a rate $\kappa_{b}/2\pi\approx 40~{}\mathrm{MHz}$ (Fig. 1a). The memory and buffer modes are coupled through a nonlinear element implementing a two-to-one photon exchange Hamiltonian $\hat{H}=\hbar g_{2}\hat{m}^{2}\hat{b}^{\dagger}+h.c.$ at a rate $g_{2}$. Here, $\hat{m}$ and $\hat{b}$ represent the annihilation operators of the memory and buffer modes, respectively. This interaction is realized in our experiment using a 3-wave mixing interaction under the condition $2\omega_{m}=\omega_{b}$, characteristic of autoparametric systems [19, 24]
Operating in the regime $\kappa_{b}\geq 8g_{2}|\alpha|$ ensures that photons in the buffer mode are evacuated into the environment at a faster rate than they are converted into pairs of photons in the memory. This allows for the adiabatic elimination of the buffer mode, leading to an effective dissipation operator $\sqrt{\kappa_{2}}\hat{m}^{2}$ where $\kappa_{2}=4g_{2}^{2}/\kappa_{b}$. In the autoparametric cat device [19], an external flux bias $\phi_{\mathrm{ext}}$ can be used to activate or deactivate two-photon dissipation. We have identified a specific flux point, $\phi_{\mathrm{on}}$, at which $\omega_{b}=2\omega_{m}$, resulting in $\kappa_{2}/2\pi=2.16~{}\mathrm{MHz}$. Another flux point, $\phi_{\mathrm{off}}$ with negligible $\kappa_{2}/2\pi$ serves to disable two-photon dissipation (details in [24]). It is chosen at a point where the self-Kerr effect is weak.
@@
Various parity measurement methods have been demonstrated in superconducting circuits, usually relying on coupling the memory at resonance [25, 26] or dispersively[27, 28, 29, 30, 31] to an auxiliary qubit. However, all these methods deteriorate as the average photon number $\bar{n}$ after displacement increases. Let us focus on the Ramsey-like sequence proposed in [27]. The qubit is first prepared in an equal superposition of ground and excited state $(|g\rangle+|e\rangle)/\sqrt{2}$ by applying an unconditional $\pi/2$ pulse, much shorter than the inverse dispersive coupling, which is $(\chi_{\mathrm{q,m}})^{-1}\approx 0.9~{}\mu\mathrm{s}$ in our experiment. After a time $T_{\mathrm{parity}}=\pi/\chi_{\mathrm{q,m}}$, the superposition accumulates a phase $n\pi$ with $n$ the number of photons of the memory. Even and odd numbers of photons are then respectively mapped onto the states (|$\left(\ket{g}+\ket{e}\right)/\sqrt{2}$ and (|$\left(\ket{g}-\ket{e}\right)/\sqrt{2}$ of the qubit. A final unconditional $\pi/2$ pulse is applied, mapping the memory parity operator $\hat{\mathcal{P}}$ onto the Pauli operator $\hat{\sigma}_{z}=|e\rangle\langle e|-|g\rangle\langle g|$ of the qubit, which can then be readout.
@@
To illustrate typical errors of Wigner tomography, we prepare a cat state |$\ket{C_{\alpha}^{+}}\propto|\alpha\rangle+|-\alpha\rangle$ and realize its tomography. The preparation is performed by tuning the flux to $\phi_{\mathrm{on}}$, which turns on the two-photon dissipation, and driving the buffer mode on resonance for 300 ns with a drive amplitude $\epsilon_{d}=-g_{2}^{*}\alpha^{2}$. As demonstrated in Refs. [9, 19], this implements an effective dissipation operator $\hat{L}_{2}=\sqrt{\kappa_{2}}\left(\hat{m}^{2}-\alpha^{2}\right)$. Fig. 2b shows the obtained Wigner tomography for $\alpha=3.3$. The typical fringes of cat states can be seen close to $\beta=0$ but their contrast is much smaller than expected for |$\ket{C_{\alpha}^{+}}$. Part of this loss of contrast comes from single photon losses during preparation. However, in the cut $W_{\hat{\rho}}(\beta)$ along $\beta\in i\mathbb{R}$ (Fig. 2.d), one can see that the measured Wigner function (orange dots) does not faithfully reproduce numerical simulations only accounting for the preparation infidelity (grey line). Here, most of the tomography errors can be attributed to single photon loss during parity measurement. Indeed, the probability of losing a photon during $T_{\mathrm{parity}}\approx 2.7~{}\mu\mathrm{s}$ reads $1-e^{-0.26\bar{n}}$ when the memory has $\bar{n}$ photons on average, reaching $50~{}\%$ for 2.7 photons. The error is therefore dramatic for the cat state in Fig. 2.b which has more than 10 photons on average. This effect is reproduced by simulations (orange line in Fig. 2.d). In contrast, the measurement of $W_{\hat{\rho}}(\pm\alpha)$ is only marginally affected by single photon loss. Indeed, after a displacement $\hat{\mathcal{D}}\left(\pm\alpha\right)$, the memory ends up in a superposition of $|0\rangle$ and |±$|\pm 2\alpha\rangle$. The parity of $|0\rangle$ does not depart from 1 during measurement and the parity of |±$|\pm 2\alpha\rangle$ is negligible, leading to a measured ${\hat{\rho}}(\pm\alpha)\approx 1/\pi$ as expected.
@@
We solve this issue by quickly turning the two-photon dissipation back on by tuning the flux to $\phi_{\mathrm{on}}$ after the displacement (Fig. 2a). Idling at $\phi_{\mathrm{on}}$ for $300~{}\mathrm{ns}$ (a few $\kappa_{2}^{-1}$) then ensures that the memory decays to a state with at most 1 photon while preserving its parity (Fig. 1b). Tuning the flux back to $\phi_{\mathrm{off}}$ enables parity measurement. In our example of a cat state above, a better measurement of the Wigner fringes can then be achieved (Fig. 2c,d). Simulations corresponding to this parity measurement protocol (blue line) predict a drastic improvement of the fringe reconstruction, reproducing what is expected when only accounting for preparation infidelity (grey line). This is confirmed by the measured Wigner function (blue dots) which shows an increase in contrast compared to the usual parity measurement protocol. It should be noted that avoiding single-photon loss while removing pairs of photons requires a large $\kappa_{2}/\kappa_{1}$ ratio. This condition can be met with other devices than the autoparametric cat, such as a simple parametrically pumped transmon qubit [11], making this method accessible to a wide range of experiments using superconducting qubits.
@@
Figure 2: (a) Pulse sequence of the dissipation enhanced Wigner tomography of |$\ket{C_{\alpha}^{+}}$ including an extra 300 ns idle time at $\phi_{\mathrm{ON}}$ after memory displacement. (b) Left: Wigner tomography of |$\ket{C_{\alpha}^{+}}$, $\alpha=3.3$, without the extra idle time. (c) Wigner tomography of |$\ket{C_{\alpha}^{+}}$ using the dissipation enhancement. (d) Cuts of the Wigner function along the imaginary axis $\mathrm{Re(\beta)}=0$ in various scenarios. Solid gray line: simulated memory state at the end of the 300 ns buffer pulse $\epsilon_{d}$. Orange points: Wigner tomography in (b). Solid orange line: simulation of the measurement protocol without the extra idle time. Blue points: Wigner tomography in (c). Solid blue line: simulation of the dissipation enhanced measurement protocol.
@@
Figure 3: (a) Inflation and deflation processes coherently map the manifolds $\mathrm{span}\left(\ket{\alpha},\ket{-\alpha}\right)$ and $\mathrm{span}\left(\ket{0},\ket{1}\right)$ by turning on or off the buffer drive $\epsilon_{d}$ [22, 18]. (b) Pulse sequence used to prepare an initial state |$\ket{\alpha}$, apply the holonomic $\hat{X}\left(\theta\right)$ gate, and measure the memory Wigner function. (c) Schematic evolution in the memory phase space of a coherent state while applying the holonomic $\hat{X}\left(\theta\right)$ gate. (d) Measured evolution of the memory Wigner function, starting from the coherent state |$\ket{-2.27}$, during the $\hat{X}\left(\pi/2\right)$ gate. The last snapshot occurs right before the final virtual rotation $\hat{R}(-\pi/2)$. An animation showing every measured time can be found as supplementary material [24].
@@
We implement a holonomic $\hat{X}\left(\theta\right)$ gate by modulating the buffer drive amplitude $\epsilon_{d}(\alpha)$ while remaining at $\phi_{\mathrm{on}}$ (Fig. 3b). We first turn off the buffer drive to deflate the memory state into the $\{|0\rangle,|1\rangle\}$ manifold. The buffer drive is then turned back on with an additional phase $2\theta$, $\epsilon_{d}\rightarrow\epsilon_{d}\left(\alpha\right)e^{2i\theta}$ [24], mapping the memory into the $\{|\alpha e^{i\theta}\rangle,|-\alpha e^{i\theta}\rangle\}$ manifold. This additional phase implements a virtual $\hat{X}(\theta)$ gate in the $\{|0\rangle,|1\rangle\}$ manifold (see Fig. 3a). A counter-rotation $\hat{R}\left(-\theta\right)$ is finally applied to bring the memory back into the cat code while preserving the phase of the superposition. This final rotation could be performed by adiabatically sweeping the phase of the buffer drive back to 00. Equivalently, we choose to apply an instantaneous virtual operation by changing the phase of the local oscillator of the memory drive by $-\theta$. The evolution of the memory in phase space starting from |$\ket{-\alpha}$ is represented in Fig. 3c. We probe the evolution of the memory during the $\hat{X}\left(\pi/2\right)$ gate starting from the coherent state |$\ket{-\alpha}$ with $\alpha=2.27$. This particular gate has been used for tomography without a transmon in [18]. Wigner tomographies of the memory are shown in Fig. 3d at $6$ different times. The memory state can first be seen evolving from the coherent state towards (|$\left(\ket{0}-\ket{1}\right)/\sqrt{2}$ (third snapshot) before being mapped into (|$\left(\ket{\alpha e^{i\pi/2}}-i\ket{-\alpha e^{i\pi/2}}\right)/\sqrt{2}$ (last snapshot).
@@
We show in Figs. 5 of [24] the state of the memory after applying the $\hat{X}\left(\theta\right)$ gate on the same coherent state |$\ket{-\alpha}$ for an arbitrary angle $\theta$. Note that the shape of the buffer drive during the gate was optimized with numerical simulations shown in Fig. S3 of [24]. The gate error being substantial, we settle for characterizing the gate properties by its effect on the coherent state $|-\alpha\rangle$ only. We compute the trace distance $T\left(\hat{\rho},|\psi(\theta)\rangle\langle\psi(\theta)|\right)$ [38, 24] between the density matrix $\hat{\rho}$ reconstructed from the measured Wigner function and the state $|\psi(\theta)\rangle=\hat{X}\left(\theta\right)|-\alpha\rangle$. The largest distance occurs at $\theta=\pm\pi/2$, where it reaches $0.23$ (see Fig. S4 in [24]). We attribute part of this apparent error to the remaining measurement infidelity (deviations between blue dots and grey line in Fig. 2d). For $\theta$ close to multiples of $\pi$, the Wigner tomography of a coherent state does not lead to such measurement errors and simulations of the memory evolution during the gate reproduce the observed trace distance. Additionally, simulations with and without the dephasing rate $\kappa_{\varphi}^{\mathrm{m}}/2\pi=0.08~{}$MHz reveal it has a major contribution to the gate error owing to the induced decoherence when the memory is in the $\{|0\rangle,|1\rangle\}$ manifold. The remaining gate errors can be attributed to a non-adiabatic evolution of the memory during the drive, or single photon loss.

### v2

@@ Abstract
Dissipation engineering offers a powerful tool for quantum technologies. Recently, new superconducting devices demonstrated an engineered two-photon dissipation rate exceeding all other relevant timescales. In particular, they have proven most useful to prevent transitions between the logical states |±$|\pm\alpha\rangle$ of a cat qubit. Here, we present three key applications of strong two-photon dissipation for quantum measurement and control, beyond cat qubit stabilization. Firstly, we demonstrate its efficacy in overcoming limitations encountered in Wigner tomography at high photon numbers. Secondly, we showcase its potential for realizing universal gates on cat qubits, exploiting the coherent mapping between cat qubit states and superpositions of 0 and 1 photons. Finally, we harness the transient dynamics of a cat state under two-photon dissipation to prepare squeezed cat states with a squeezing factor exceeding $3.96\pm 0.07$ dB.
I Introduction
@@
To implement such a two-photon dissipation, a dedicated mode, denoted as the buffer, is used to mediate the interaction between the memory and its environment [9, 11, 12, 16, 18, 19]. In our experiment, this buffer dissipates photons at a rate $\kappa_{b}/2\pi\approx 40~{}\mathrm{MHz}$ (Fig. 1a). The memory and buffer modes are coupled through a nonlinear element implementing a two-to-one photon exchange Hamiltonian $\hat{H}=\hbar g_{2}\hat{m}^{2}\hat{b}^{\dagger}+h.c.$ at a rate $g_{2}$. Here, $\hat{m}$ and $\hat{b}$ represent the annihilation operators of the memory and buffer modes, respectively. This interaction is realized in our experiment using a 3-wave mixing interaction under the condition $2\omega_{m}=\omega_{b}$, characteristic of autoparametric systems [19] (see Sec. A).
Operating in the regime $\kappa_{b}\geq 8g_{2}|\alpha|$ ensures that photons in the buffer mode are evacuated into the environment at a faster rate than they are converted into pairs of photons in the memory. This allows for the adiabatic elimination of the buffer mode, leading to an effective dissipation operator $\sqrt{\kappa_{2}}\hat{m}^{2}$ where $\kappa_{2}=4g_{2}^{2}/\kappa_{b}$. In the autoparametric cat device [19], an external flux bias $\phi_{\mathrm{ext}}$ can be used to activate or deactivate two-photon dissipation. We have identified a specific flux point, $\phi_{\mathrm{on}}$, at which $\omega_{b}=2\omega_{m}$, resulting in $\kappa_{2}/2\pi=2.16~{}\mathrm{MHz}$. Another flux point, $\phi_{\mathrm{off}}$ with negligible $\kappa_{2}/2\pi$ serves to disable two-photon dissipation (details in Sec. B.1). It is chosen at a point where the self-Kerr effect is weak.
@@ II Dissipation enhanced Wigner tomography
Various parity measurement methods have been demonstrated in superconducting circuits, usually relying on coupling the memory at resonance [24, 25] or dispersively [26, 27, 28, 29, 30] to an auxiliary qubit. However, all these methods deteriorate as the average photon number $\bar{n}$ after displacement increases. Let us focus on the Ramsey-like sequence proposed in Ref. [26]. The qubit is first prepared in an equal superposition of ground and excited state $(|g\rangle+|e\rangle)/\sqrt{2}$ by applying an unconditional $\pi/2$ pulse, much shorter than the inverse dispersive coupling, which is $(\chi_{\mathrm{q,m}})^{-1}\approx 0.9~{}\mu\mathrm{s}$ in our experiment. After a time $T_{\mathrm{parity}}=\pi/\chi_{\mathrm{q,m}}$, the superposition accumulates a phase $n\pi$ with $n$ the number of photons of the memory. Even and odd numbers of photons are then respectively mapped onto the states (|$\left(\ket{g}+\ket{e}\right)/\sqrt{2}$ and (|$\left(\ket{g}-\ket{e}\right)/\sqrt{2}$ of the qubit. A final unconditional $\pi/2$ pulse is applied, mapping the memory parity operator $\hat{\mathcal{P}}$ onto the Pauli operator $\hat{\sigma}_{z}=|e\rangle\langle e|-|g\rangle\langle g|$ of the qubit, which can then be readout.
@@
To illustrate typical errors of Wigner tomography, we prepare a cat state |$\ket{C_{\alpha}^{+}}\propto|\alpha\rangle+|-\alpha\rangle$ and realize its tomography. The preparation is performed by tuning the flux to $\phi_{\mathrm{on}}$, which turns on the two-photon dissipation, and driving the buffer mode on resonance for 300 ns with a drive amplitude $\epsilon_{d}=-g_{2}^{*}\alpha^{2}$. As demonstrated in Refs. [9, 19], this implements an effective dissipation operator $\hat{L}_{2}=\sqrt{\kappa_{2}}\left(\hat{m}^{2}-\alpha^{2}\right)$. Fig. 2b shows the obtained Wigner tomography for $\alpha=3.3$. The typical fringes of cat states can be seen close to $\beta=0$ but their contrast is much smaller than expected for |$\ket{C_{\alpha}^{+}}$. Part of this loss of contrast comes from single photon losses during preparation. However, in the cut $W_{\hat{\rho}}(\beta)$ along $\beta\in i\mathbb{R}$ (Fig. 2.d), one can see that the measured Wigner function (orange dots) does not faithfully reproduce numerical simulations only accounting for the preparation infidelity (grey line). Here, most of the tomography errors can be attributed to single photon loss during parity measurement. Indeed, the probability of losing a photon during $T_{\mathrm{parity}}\approx 2.7~{}\mu\mathrm{s}$ reads $1-e^{-0.26\bar{n}}$ when the memory has $\bar{n}$ photons on average, reaching $50~{}\%$ for 2.7 photons. The error is therefore dramatic for the cat state in Fig. 2.b which has more than 10 photons on average. This effect is reproduced by simulations (orange line in Fig. 2.d). In contrast, the measurement of $W_{\hat{\rho}}(\pm\alpha)$ is only marginally affected by single photon loss. Indeed, after a displacement $\hat{\mathcal{D}}\left(\pm\alpha\right)$, the memory ends up in a superposition of $|0\rangle$ and |±$|\pm 2\alpha\rangle$. The parity of $|0\rangle$ does not depart from 1 during measurement and the parity of |±$|\pm 2\alpha\rangle$ is negligible, leading to a measured $W_{\hat{\rho}}(\pm\alpha)\approx 1/\pi$ as expected (see Fig. 6).
@@
We solve this issue by quickly turning the two-photon dissipation back on by tuning the flux to $\phi_{\mathrm{on}}$ after the displacement (Fig. 2a). Idling at $\phi_{\mathrm{on}}$ for $300~{}\mathrm{ns}$ (a few $\kappa_{2}^{-1}$) then ensures that the memory decays to a state with at most 1 photon while preserving its parity (Fig. 1b). Tuning the flux back to $\phi_{\mathrm{off}}$ enables parity measurement. In our example of a cat state above, a better measurement of the Wigner fringes can then be achieved (Fig. 2c,d). Simulations corresponding to this parity measurement protocol (blue line) predict a drastic improvement of the fringe reconstruction, reproducing up to $5~{}\%$ what is expected when only accounting for preparation infidelity (grey line). This is confirmed by the measured Wigner function (blue dots) which shows an increase in contrast compared to the usual parity measurement protocol. It should be noted that avoiding single-photon loss while removing pairs of photons requires a larger $\kappa_{2}/\kappa_{1}$ ratio. This condition can be met with other devices than the autoparametric cat, such as a simple parametrically pumped transmon qubit [11], making this method accessible to a wide range of experiments using superconducting qubits. Note that we do not find an upper bound in photon number for this tomography technique and manage to demonstrate it up to 25 photons. We note that this protocol was used in Ref. [19].
@@
Figure 2: (a) Pulse sequence of the dissipation enhanced Wigner tomography of |$\ket{C_{\alpha}^{+}}$ including an extra 300 ns idle time at $\phi_{\mathrm{ON}}$ after memory displacement. (b) Left: Wigner tomography of |$\ket{C_{\alpha}^{+}}$, $\alpha=3.3$, without the extra idle time. (c) Wigner tomography of |$\ket{C_{\alpha}^{+}}$ using the dissipation enhancement. (d) Cuts of the Wigner function along the imaginary axis $\mathrm{Re(\beta)}=0$ in various scenarios. Solid gray line: simulated memory state at the end of the 300 ns buffer pulse $\epsilon_{d}$. Orange points: Wigner tomography in (b). Solid orange line: simulation of the measurement protocol without the extra idle time. Blue points: Wigner tomography in (c). Solid blue line: simulation of the dissipation enhanced measurement protocol (see Sec. B.3).
III Quantum gates on cat qubits
@@
Figure 3: (a) Inflation and deflation processes coherently map the manifolds $\mathrm{span}\left(\ket{\alpha},\ket{-\alpha}\right)$ and $\mathrm{span}\left(\ket{0},\ket{1}\right)$ by turning on or off the buffer drive $\epsilon_{d}$ [22, 18]. (b) Pulse sequence used to prepare an initial state |$\ket{\alpha}$, apply the holonomic $\hat{X}\left(\theta\right)$ gate, and measure the memory Wigner function. (c) Schematic evolution in the memory phase space of a coherent state while applying the holonomic $\hat{X}\left(\theta\right)$ gate. (d) Measured evolution of the memory Wigner function, starting from the coherent state |$\ket{-2.27}$, during the $\hat{X}\left(\pi/2\right)$ gate. The last snapshot occurs right before the final virtual rotation $\hat{R}(-\pi/2)$. Every measured time can be found in the appendix Video C.2.3.
@@
We implement a holonomic $\hat{X}\left(\theta\right)$ gate by modulating the buffer drive amplitude $\epsilon_{d}(\alpha)$ while remaining at $\phi_{\mathrm{on}}$ (Fig. 3b). We first turn off the buffer drive to deflate the memory state into the $\{|0\rangle,|1\rangle\}$ manifold. The buffer drive is then turned back on with an additional phase $2\theta$, $\epsilon_{d}\rightarrow\epsilon_{d}\left(\alpha\right)e^{2i\theta}$ (see Sec. C.2), mapping the memory into the $\{|\alpha e^{i\theta}\rangle,|-\alpha e^{i\theta}\rangle\}$ manifold. This additional phase implements a virtual $\hat{X}(\theta)$ gate in the $\{|0\rangle,|1\rangle\}$ manifold (see Fig. 3a). A rotation in the memory phase space $\hat{R}\left(-\theta\right)$, which maps any coherent state $|\beta e^{i\theta}\rangle$ onto $|\beta\rangle$, is finally applied to bring the memory back into the cat code while preserving the phase of the superposition. This final rotation could be performed by adiabatically sweeping the phase of the buffer drive back to 00. Equivalently, we choose to apply an instantaneous virtual operation by changing the phase of the local oscillator of the memory drive by $-\theta$. The evolution of the memory in phase space starting from |$\ket{-\alpha}$ is represented in Fig. 3c. We probe the evolution of the memory during the $\hat{X}\left(\pi/2\right)$ gate starting from the coherent state |$\ket{-\alpha}$ with $\alpha=2.27$. This particular gate has been used for tomography without a transmon in [18]. Wigner tomographies of the memory are shown in Fig. 3d at $6$ different times. The memory state can first be seen evolving from the coherent state towards (|$\left(\ket{0}-\ket{1}\right)/\sqrt{2}$ (third snapshot) before being mapped into (|$\left(\ket{\alpha e^{i\pi/2}}-i\ket{-\alpha e^{i\pi/2}}\right)/\sqrt{2}$ (last snapshot).
@@
We show in Fig. 10 and Video C.2.2 the state of the memory after applying the $\hat{X}\left(\theta\right)$ gate on the same coherent state |$\ket{-\alpha}$ for an arbitrary angle $\theta$. Note that the shape of the buffer drive during the gate was optimized with numerical simulations shown in Fig. 8. The gate error being substantial, we settle for characterizing the gate properties by its effect on the coherent state $|-\alpha\rangle$ only. We compute the trace distance $T\left(\hat{\rho},|\psi(\theta)\rangle\langle\psi(\theta)|\right)$ [37] (see Eq. (8)) between the density matrix $\hat{\rho}$ reconstructed from the measured Wigner function and the state $|\psi(\theta)\rangle=\hat{X}\left(\theta\right)|-\alpha\rangle$. The largest distance occurs at $\theta=\pm\pi/2$, where it reaches $0.23$ (see Fig. 10). Numerical simulations qualitatively reproduce the dependence of the trace distance as a function of $\theta$. Larger trace distances are obtained for angles close to $\theta=\pm\pi/2$, corresponding to states exhibiting coherence, owing to the impact of phase flip errors during the inflation and stabilization step of the gate. Additionally, deviation between the simulations and measurements can be observed at those angles, which we explain by measurement infidelity (deviations between blue dots and grey line in Fig. 2d). Note that measurement infidelity is larger for cat states than for coherent states. Finally, simulations with and without the dephasing rate $\kappa_{\varphi}^{\mathrm{m}}/2\pi=0.08~{}$MHz reveal it has a major contribution to the gate error owing to induced decoherence when the memory is in the $\{|0\rangle,|1\rangle\}$ manifold. The remaining gate errors observed when $\kappa_{\varphi}^{\mathrm{m}}/2\pi=0.00~{}$MHz (blue line in Fig. 10b) can be attributed to a non-adiabatic evolution of the memory during the gate, or single photon loss.
IV Preparing squeezed cat states